OpenStudy (anonymous):

please help! The bacteria in a certain culture double every 7.3 hours. The culture has 6,500 bacteria at the start. How many bacteria will the culture contain after 5 hours

OpenStudy (anonymous):

we can start with \(6,500\times 2^{\frac{t}{7.3}}\)

OpenStudy (anonymous):

1805.56t

OpenStudy (anonymous):

could be

OpenStudy (anonymous):

did you try \[65,000\times2^{\frac{5}{7.3}}\]?

OpenStudy (anonymous):

89041.1

OpenStudy (anonymous):

is there more steps? none of my answer choices relate to that it seems better than what i did before though for sure

OpenStudy (anonymous):

this is what i get http://www.wolframalpha.com/input/?i=65000*2^%285%2F7.3%29